OpenStudy (anonymous):

If f(x) = x2 - 1, find f(3).

OpenStudy (anonymous):

whenever you see an x replace it with a 3 to get your answer. f(3)=3^2-1=9-1=8

OpenStudy (anonymous):

yeah. correct! @luis and zed